calculate the volume of each cone. use 3.14 and round eachyour answer to the nearest hundreatheth if neccesery
a traffic cone has a radius of 9in the height of the cone is 30

Answer by checkley77(12844)

V=BH/3
V=(PIR^2*30)/3
V=3.14*9^2*30/3
V=3.14*81*10
V=2,543.4 IN^3